# Routing Guidance for Emerging Transportation Systems with Improved Dynamic Trip Equity

### v2

@@ Abstract
In this paper, we present a dynamic routing guidance system that optimizes route recommendations for individual vehicles within an emerging transportation system while enhancing travelers’ trip equity. We develop a framework to quantify trip quality and equity in a dynamic travel environment, providing new insights into how routing guidance influences equity in road transportation. Our approach enables real-time routing by incorporating both monitored and anticipated traffic congestion. We provide conditions that ensure achieving perfect trip equity for all travelers in a free-flow network. Finally, simulation studies on 1,000 vehicles traversing an urban road network in Boston demonstrate that our proposed method improves trip equity by approximately 11.4% compared to the shortest-route strategy. In addition, the results reveal that our approach redistributes travel costs across vehicle types through route optimization, contributing to a more equitable transportation system.
@@ I Introduction
Equity has been an important principle in transportation to promote fairness and inclusivity [5]. It advocates for the just distribution of transportation resources, services, and infrastructure, ensuring that all individuals, regardless of their socioeconomic status, race, gender, or geographic location, have equal access to safe, affordable, and reliable mobility options. Transportation solutions lacking access equity can lead to unintended negative consequences [6]. As noted in [7], low-income populations, who often rely on public transit, face substantial barriers when affordable and reliable transportation options are limited. This inequity restricts their access to essential services and opportunities and reinforces cycles of poverty and social exclusion. In addition, disparities in transportation access can drive over-reliance on private vehicles, exacerbating traffic congestion, increasing travel costs, and contributing to environmental degradation.
Advancing equity in emerging transportation systems requires developing routing guidance systems that address the diverse needs of travelers, such as availability and affordability, while prioritizing equity. Different from conventional routing guidance services [8], which focus on optimizing traffic efficiency through metrics such as shortest distance, minimal time, or cost, equity-based routing guidance systems integrate fairness into existing frameworks, delivering socially optimal solutions for every traveler despite their disparities.
@@
In this paper, we present a D-RGS designed to optimize the route recommendations for individual vehicles in an emerging transportation system with enhanced trip equity. Unlike existing D-RGSs, our approach integrates equity considerations into the decision-making process, ensuring that the benefits of transportation advancements are distributed fairly across diverse populations. By leveraging emerging technologies, including automated and ride-hailing vehicles, our system aims to bridge the gap between socioeconomic groups, reduce racial and ethnic disparities, and promote sustainable transportation alternatives. The contributions of this work are summarized as follows:
  • We present a novel framework to measure the quality and equity of trips in a dynamic emerging transportation system through the notions of the dynamic trip index and dynamic trip equity.
  • Based on the proposed metrics, we derive conditions for achieving perfect trip equity in free-flow networks, which provides the upper bound for trip equity attainable through route optimization.
  • We propose a D-RGS that incorporates both monitored and estimated traffic congestion in real-time route optimization and allows for trip equity improvement.
Finally, simulation studies on $1,000$ vehicles traversing an urban road network in Boston demonstrate the effectiveness of the proposed framework.
@@ II Problem Formulation
In this section, we formulate the routing problem for individual vehicles. We start by introducing the road network, vehicle dynamics, and traffic flow. Subsequently, we formally present the routing problem.
@@ II-A Road Network
We model the road network as a directed graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes and $\mathcal{E}\!\subset\!{\mathcal{V}\!\times\!{\mathcal{V}}}$ is the set of edges. Each node $v\!\in\!\mathcal{V}$ is uniquely indexed, representing a key intersection or an origin/destination node in the network, while each edge $(v,v^{\prime})\!\in\!\mathcal{E}$ represents a road segment connecting a pair of nodes. The weight of each edge, denoted by $\tau(v,v^{\prime})\!\in\!\mathbb{R}_{+}$, denotes the travel time to traverse the edge, where $\mathbb{R}_{+}$ is the set of positive real numbers. Throughout this paper, we assume all vehicles travel with a constant free-flow speed. We denote the nominal (i.e., free-flow) travel time on an edge $(v,v^{\prime})$ by $\tau_{0}(v,v^{\prime})$.
Consider an emerging transportation system consisting of diverse transportation modes, including private vehicles, autonomous vehicles, and ride-hailing vehicles. At any time $t$, let $N(t)\!\in\!{\mathbb{N}}$ be the number of vehicles actively traveling in the network. By “actively,” we refer to the vehicle taking at least one traveler en route to a designated destination. For any vehicle $i\!\in\!\mathcal{N}(t):=\{1,2,\dots,N(t)\}$ traveling from a node $v_{1}$ to node $v_{n}$, a feasible route between the two nodes is denoted as
@@ II-B Vehicle Dynamics
To describe vehicle movements between DMPs, we introduce the following dynamic model. For each vehicle $i\!\in\!\mathcal{N}(t)$, let $a_{k}^{i}$ and $a_{k+1}^{i}$ be its arrival times at two consecutive DMPs $v_{k}$ and $v_{k+1}$, respectively. The dynamics of vehicle $i$ is then depicted as
@@
where $\tau_{k,k+1}^{i}$ denotes the travel time of vehicle $i$ to traverse the edge $(v_{k},v_{k+1})$, which can be modeled using the widely adopted Bureau of Public Road (BPR) function [13]. The BPR model captures the impact of traffic congestion by describing how travel time increases as traffic demand approaches or exceeds road capacity. Specifically, $\tau_{k,k+1}^{i}$ can be denoted as
@@
In a dynamic system, $f_{k,k+1}^{i}$ varies over time, depending on the number of vehicles traveling on $(v_{k},v_{k+1})$ within the same time interval. Next, we present appropriate monitoring and estimation approaches to characterize the traffic flow for each vehicle based on its arrival time at a DMP.
II-C Traffic Flow Monitoring &\&& Estimation
@@ II-C1 Flow Monitoring for Adjacent Edges
$\mathcal{F}\big{(}a_{k}^{i},\Delta{t}\big{)}\!:=\!\big{[}a_{k}^{i}\!-\!\Delta{t},a_{k}^{i}\!+\!\Delta{t}\big{]},$ (4)
where $\Delta{t}\!>\!{0}$ defines the size of the monitoring window. Accounting for other vehicles $j\!\in\!\mathcal{N}(t)\!\setminus\!\{i\}$ entering the same edge within this window, we define the following indicator function
$\!\!\mathbf{1}_{a_{k,k+1}^{j}}\!:=\!\begin{cases}1,&\text{if~{} $a_{k,k+1}^{j}\!\in\!{\mathcal{F}\big{(}a_{k}^{i},\Delta{t}\big{)}}$,}\\ 0,&\text{otherwise,}\end{cases}$ (5)
@@ II-C2 Flow Estimation for Future Edges
For future edges $(v_{s},v_{s+1})\!\in\!{r_{k,d_{i}}}$, $s\!\neq\!{k}$, the traffic flow at a future arrival time $a_{s}^{i}$ cannot be directly observed at time $a_{k}^{i}$. However, by leveraging the route plans of other vehicles $j\!\in\!\mathcal{N}(t)\!\setminus\!\{i\}$, the traffic flow can be anticipated, enabling congestion estimation and route optimization for vehicle $i$ at each of its DMP $v_{k}$.
@@
where $\tilde{\tau}_{k,k+1}^{i}$ representing the time for vehicle $i$ to traverse edge $(v_{k},v_{k+1})$ is computed by integrating the monitored traffic flow $\tilde{f}_{k,k+1}^{i}$ in (3). Additionally, $\hat{\tau}_{s-1,s}^{i}$ in (7) denotes the estimated travel time for vehicle $i$ to traverse the future edge $(v_{s-1},v_{s})$. Using the estimated arrival time $\hat{a}_{s}^{i}$, the traffic flow on the future edge $(v_{s},v_{s+1})$ can be estimated by
@@
$\mathbf{1}_{\hat{a}_{s,s+1}^{j}}\!:=\!\begin{cases}1,&\text{if~{}$\hat{a}_{s,s+1}^{j}\!\in\!{\mathcal{F}\big{(}\hat{a}_{s}^{i},\Delta{t}\big{)}}$,}\\ 0,&\text{otherwise.}\end{cases}$ (9)
By integrating $\hat{f}_{s,s+1}^{i}$ into (3), the estimated travel time $\hat{\tau}_{s,s+1}^{i}$ and the estimated arrival time $\hat{a}_{s+1}^{i}$ can be obtained accordingly. Note that for each future edge $(v_{s},v_{s+1})\!\in\!{r_{k,d_{i}}}$, $s\!\neq\!{k}$, the estimated traffic flow $\hat{f}_{s,s+1}^{i}$, travel time $\hat{\tau}_{s,s+1}^{i}$, and the resulting estimated arrival time $\hat{a}_{s+1}^{i}$ are computed iteratively.
@@ II-D Routing Problem
In our exposition, we consider an emerging transportation system that integrates diverse transportation modes, including (1) private vehicles, (2) autonomous vehicles, and (3) ride-hailing vehicles. The sets of vehicles in each category are denoted by $\mathcal{N}_{p}(t)$, $\mathcal{N}_{a}(t)$, and $\mathcal{N}_{h}(t)$, respectively, with $\mathcal{N}(t)\!=\!\mathcal{N}_{p}(t)\cup{\mathcal{N}_{a}(t)}\cup{\mathcal{N}_{h}(t)}$. For simplicity, we assume travelers taking ride-hailing vehicles share the same OD pairs and evenly split the travel expenses. Private and autonomous vehicles, however, serve only one traveler at a time. Owing to the automation technology, the per-mile travel cost of autonomous vehicles is typically lower than that of private vehicles.
While individuals select their travel modes based on factors such as income, accessibility, and reliability, the quality of their trips should remain comparable to ensure equitable access to transportation resources and services. To this end, this paper aims to develop a D-RGS that assists each vehicle $i\!\in\!\mathcal{N}(t)$ in route planning by optimally selecting a route $r_{k,d_{i}}^{*}$ from $\mathcal{R}_{k,d_{i}}$ at each DMP $v_{k}$ while enhancing trip equity across all travelers.
@@ III Dynamic Trip Index and Trip Equity
This section presents a framework to evaluate the quality of each trip and the equity of all trips in a dynamic transportation system. The framework is built upon two concepts: the dynamic trip index and dynamic trip equity that we introduce next.
@@ III-A Dynamic Trip Index
$T_{w,\min}^{i}:\!$ $=\min\big{\{}T_{w,p},T_{w,a},T_{w,h}\big{\}},$ (12a)
$T_{d,\max}^{i}:\!$ $=\max\big{\{}T_{d,p},T_{d,a},T_{d,h}\big{\}}.$ (12b)
We make use of $T_{w,p}$, $T_{w,a}$, $T_{w,h}$ to denote the average waiting times to access a private, autonomous, and ride-hailing vehicle, respectively. Here, $T_{w,p}$ accounts for the additional time required to park and retrieve a private vehicle, $T_{w,a}$ captures the potential delays due to scheduling and operational constraints of an autonomous vehicle, and $T_{w,h}$ reflects the average waiting time to dispatch a ride-hailing vehicle. In (12b), $T_{d,p}$, $T_{d,a}$, $T_{d,h}$ denote the feasible departure time windows (in hours per day) for each type of vehicle. It is worth noting that $T_{w,i}$ quantifies the accessibility of vehicle $i$, while $T_{d,i}$ characterizes its availability.
@@
Remark 1
The values of $\tau_{\min}^{i}$, $\phi_{\min}^{i}$, and $q_{\min}^{i}$ are achieved by selecting the most favorable vehicle type, which offers the maximum convenience, the shortest travel time, and operates under ideal traffic conditions without congestion. By comparing the actual travel conditions to this optimal scenario, the dynamic metric $DTX_{i}(t)$ quantifies the trip quality of the traveler in vehicle $i$ relative to the best possible travel experience.
@@ III-B Dynamic Trip Equity
Instead of evaluating the DTE among all travelers in $\mathcal{N}(t)$, which can be computationally inefficient, we propose the DTE defined within the RRCs for each vehicle based on their route plans, ignoring indirect couplings between vehicles. Let $\mathcal{C}_{i,p}(t)$, $\mathcal{C}_{i,a}(t)$, $\mathcal{C}_{i,h}(t)$ be the sets of private, autonomous, and ride-hailing vehicles in $\mathcal{C}_{i}(t)$, i.e., $\mathcal{C}_{i}(t)\!=\!\mathcal{C}_{i,p}(t)\cup{\mathcal{C}_{i,a}(t)}\cup{\mathcal{C}_{i,h}(t)}$. We employ the Gini Coefficient [14], an effective measure of inequity in economics and social sciences, to quantify the DTX distribution among travelers. The Gini Coefficient is built upon the Lorenz curve [15], and by its definition, the DTE for the traveler in vehicle $i$ and those in the RRCs of vehicle $i$ at $t$ is defined as
@@ III-C Dynamic Trip Equity in Free-Flow Networks
This subsection discusses the DTE in a free-flow road network. In scenarios where the infrastructure is sufficient or the number of vehicles in the system is well below the road capacity, congestion is not a concern in route planning. In these cases, vehicles can freely choose the shortest routes for their trips. To describe the trip quality in this case, let $DTX_{i,0}$ be the DTX for vehicle $i\!\in\!\mathcal{N}(t)$ traversing a free-flow network. Since $\tau_{i}(t)\!=\!\tau_{\min}^{i}$, by (10), (11b), and (14b), we have
@@
For any vehicle $i\!\in\!\mathcal{N}(t)$ traveling in a free-flow road network on the trip $(o_{i},d_{i})$, the best DTE can be achieved between the traveler in vehicle $i$ and those in its RRCs, denoted as $\mathcal{C}_{i,0}(t)$, if the following condition is satisfied:
@@
Proposition 1 reveals the conditions required for maximizing trip equity across travelers in a free-flow road network. These conditions impose constraints on road infrastructure, transportation costs, and vehicle scheduling, which collectively determine the achievable DTX for each vehicle. Note that meeting these conditions ensures that the upper bound of the DTE attainable through route optimization is $1$. In other words, the D-RGS improves trip equity only through balancing the DTX via reducing traffic congestion. However, it cannot mitigate inequities caused by other factors, such as travel costs or limitations inherent to certain vehicle types.

### v3

@@ Abstract
This paper presents a dynamic routing guidance system that optimizes route recommendations for individual vehicles in an emerging transportation system while enhancing travelers’ trip equity. We develop a framework to quantify trip quality and equity in dynamic travel environments, providing new insights into how routing guidance influences equity in road transportation. Our approach enables real-time routing by incorporating both monitored and anticipated traffic congestion. We provide conditions that ensure perfect trip equity for all travelers in a free-flow network. Simulation studies on 1,000 vehicles traversing an urban road network in Boston demonstrate that our method improves trip equity by approximately 11.4% compared to the shortest-route strategy. In addition, the results reveal that our approach redistributes travel costs across vehicle types through route optimization, contributing to a more equitable transportation system.
@@ I Introduction
Equity has been an important principle in transportation to promote fairness and inclusivity [5]. It advocates for the just distribution of transportation resources, services, and infrastructure, ensuring that all individuals, regardless of their socioeconomic status, race, gender, or geographic location, have equal access to safe, affordable, and reliable mobility options. Transportation solutions lacking access equity can lead to unintended negative consequences [6]. As noted in [7], low-income populations, who often rely on public transit, face substantial barriers when affordable and reliable transportation options are limited. This inequity restricts their access to essential services and opportunities, and reinforces cycles of poverty and social exclusion. In addition, disparities in transportation access can drive over-reliance on private vehicles, exacerbating traffic congestion, increasing travel costs, and contributing to environmental degradation.
Advancing equity in emerging transportation systems requires developing routing guidance systems that address the diverse needs of travelers, such as availability and affordability, while prioritizing equity. Different from conventional routing guidance services [8], which focus on optimizing traffic efficiency through metrics such as shortest distance, minimal time, or cost, equity-based routing guidance systems integrate fairness into existing frameworks, delivering socially optimal solutions for travelers despite their disparities.
@@
In this paper, we present a D-RGS designed to optimize the route recommendations for individual vehicles in an emerging transportation system with enhanced trip equity. Unlike existing D-RGSs, our approach integrates equity considerations into the decision-making process, ensuring that the benefits of transportation advancements are distributed fairly across diverse populations. The main contributions are as follows: (i) We present a novel framework to measure the quality and equity of trips in a dynamic emerging transportation system through the notions of the dynamic trip index and dynamic trip equity. (ii) Based on the proposed metrics, we derive conditions for achieving perfect trip equity in free-flow networks, which provide the upper bound for trip equity attainable through route optimization. (iii) We propose a D-RGS that incorporates both monitored and estimated traffic congestion in real-time route optimization and allows for trip equity improvement. Finally, simulation studies on $1,000$ vehicles traversing an urban road network in Boston demonstrate the effectiveness of the proposed framework.
@@ II-A Road Network
We model the road network as a directed graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes and $\mathcal{E}\!\subset\!{\mathcal{V}\!\times\!{\mathcal{V}}}$ is the set of edges. Each node $v\!\in\!\mathcal{V}$ is uniquely indexed, representing a key intersection or an origin/destination node in the network, while each edge $(v,v^{\prime})\!\in\!\mathcal{E}$ represents a road segment connecting a pair of nodes. The weight of each edge, denoted by $\tau(v,v^{\prime})\!\in\!\mathbb{R}_{+}$, denotes the travel time to traverse the edge, where $\mathbb{R}_{+}$ is the set of positive real numbers. We denote the free-flow travel time on an edge $(v,v^{\prime})$ by $\tau_{0}(v,v^{\prime})$.
Consider an emerging transportation system consisting of diverse transport modes, such as private vehicles and autonomous vehicles. At any time $t$, let $N(t)\!\in\!{\mathbb{N}}$ denote the number of vehicles actively traveling in the network. By “actively,” we refer to the vehicle taking at least one traveler en route to a designated destination. For vehicle $i\!\in\!\mathcal{N}(t)\!:=\!\{1,2,\dots,N(t)\}$ traveling from a node $v_{1}$ to node $v_{n}$, a feasible route between $v_{1}$ and $v_{n}$ is defined as
@@ II-B Vehicle Dynamics
To describe vehicle movements between DMPs, we introduce the following dynamic model. For each vehicle $i\!\in\!\mathcal{N}(t)$, let $a_{k}^{i}$ and $a_{k+1}^{i}$ be its arrival times at two consecutive DMPs $v_{k}$ and $v_{k+1}$, respectively. The dynamics of vehicle $i$ is depicted as
@@
where $\tau_{k,k+1}^{i}$ is the travel time of vehicle $i$ to traverse edge $(v_{k},v_{k+1})$, which can be modeled using the widely adopted Bureau of Public Road (BPR) function [13]. The BPR model captures the impact of traffic congestion by describing how travel time increases as traffic demand approaches or exceeds road capacity. Specifically, $\tau_{k,k+1}^{i}$ is of the form
@@
In a dynamic system, $f_{k,k+1}^{i}$ varies over time, depending on the number of vehicles traveling on $(v_{k},v_{k+1})$ within the same time interval. Next, we present appropriate monitoring and estimation approaches to characterize the traffic flow.
II-C Traffic Flow Monitoring &\& Estimation
@@ II-C1 Flow Monitoring for Adjacent Edges
$\mathcal{M}\big{(}a_{k}^{i},\Delta{t}\big{)}\!:=\!\big{[}a_{k}^{i}\!-\!\Delta{t},a_{k}^{i}\!+\!\Delta{t}\big{]},$ (4)
where $\Delta{t}\!>\!{0}$ denotes the size of the monitoring window. Accounting for other vehicles $j\!\in\!\mathcal{N}(t)\!\setminus\!\{i\}$ entering the same edge within this window, we define the indicator function
$\!\!\mathbf{1}_{a_{k,k+1}^{j}}\!:=\!\begin{cases}1,&\text{if~ $a_{k,k+1}^{j}\!\in\!{\mathcal{M}\big{(}a_{k}^{i},\Delta{t}\big{)}}$,}\\ 0,&\text{otherwise,}\end{cases}$ (5)
@@ II-C2 Flow Estimation for Future Edges
For future edges $(v_{s},v_{s+1})\!\in\!{r_{k,d_{i}}}$, $s\!\neq\!{k}$, the traffic flow at arrival time $a_{s}^{i}$ cannot be directly observed at time $a_{k}^{i}$. However, by leveraging the route plans of other vehicles $j\!\in\!\mathcal{N}(t)\!\setminus\!\{i\}$, the traffic flow can be anticipated, enabling congestion estimation and route optimization for vehicle $i$ at each of its DMPs.
@@
where $\tilde{\tau}_{k,k+1}^{i}$ representing the time for vehicle $i$ to traverse edge $(v_{k},v_{k+1})$ is computed by integrating the monitored traffic flow $\tilde{f}_{k,k+1}^{i}$ in (3). Additionally, $\hat{\tau}_{s-1,s}^{i}$ in (7) denotes the estimated travel time for vehicle $i$ to traverse the edge $(v_{s-1},v_{s})$. Using the estimated arrival time $\hat{a}_{s}^{i}$, the traffic flow on the future edge $(v_{s},v_{s+1})$ can be estimated by
@@
$\mathbf{1}_{\hat{a}_{s,s+1}^{j}}\!:=\!\begin{cases}1,&\text{if~ $\hat{a}_{s,s+1}^{j}\!\in\!{\mathcal{M}\big{(}\hat{a}_{s}^{i},\Delta{t}\big{)}}$,}\\ 0,&\text{otherwise.}\end{cases}$ (9)
Integrating $\hat{f}_{s,s+1}^{i}$ into (3), the estimated travel time $\hat{\tau}_{s,s+1}^{i}$ and arrival time $\hat{a}_{s+1}^{i}$ can be obtained accordingly. Note that for each future edge $(v_{s},v_{s+1})\!\in\!{r_{k,d_{i}}}$, $s\!\neq\!{k}$, the estimated traffic flow $\hat{f}_{s,s+1}^{i}$, travel time $\hat{\tau}_{s,s+1}^{i}$, and the resulting estimated arrival time $\hat{a}_{s+1}^{i}$ are computed iteratively.
@@ II-D Routing Problem
In our exposition, we consider an emerging transportation system integrating (1) private vehicles, (2) autonomous vehicles, and (3) ride-hailing vehicles. The sets of vehicles in each category are denoted by $\mathcal{N}_{p}(t)$, $\mathcal{N}_{a}(t)$, $\mathcal{N}_{h}(t)$, with $\mathcal{N}(t)\!=\!\mathcal{N}_{p}(t)\!\cup\!{\mathcal{N}_{a}(t)}\!\cup\!{\mathcal{N}_{h}(t)}$. For simplicity, we assume travelers taking ride-hailing vehicles share the same OD pairs and split costs equally, while private and autonomous vehicles serve only one traveler at a time. Owing to automation technology, autonomous vehicles have a lower per-mile cost than that of private vehicles.
While individuals select their travel modes based on factors such as income, accessibility, and reliability, the quality of their trips should remain comparable to ensure equitable access to transportation resources and services. This paper aims to develop a D-RGS guide each vehicle $i\!\in\!\mathcal{N}(t)$ in selecting an optimal route $r_{k,d_{i}}^{*}\!\in\!\mathcal{R}_{k,d_{i}}$ at each DMP $v_{k}$ while enhancing trip equity across all travelers.
@@ III Dynamic Trip Index and Trip Equity
This section presents a framework to evaluate the quality of each trip and the equity of all trips in a dynamic transportation system. The framework is built upon two key concepts: the dynamic trip index and dynamic trip equity that we introduce next.
@@ III-A Dynamic Trip Index
$T_{w,\min}^{i}\!$ $=\min\big{\{}T_{w,p},T_{w,a},T_{w,h}\big{\}},$ (12a)
$T_{d,\max}^{i}\!$ $=\max\big{\{}T_{d,p},T_{d,a},T_{d,h}\big{\}}.$ (12b)
We use $T_{w,p}$, $T_{w,a}$, $T_{w,h}$ to denote the average waiting times to access a private, autonomous, and ride-hailing vehicle, respectively. Here, $T_{w,p}$ accounts for the additional time required to park and retrieve a private vehicle, $T_{w,a}$ captures the potential delays due to scheduling and operational constraints of an autonomous vehicle, and $T_{w,h}$ reflects the average waiting time to dispatch a ride-hailing vehicle. In (12b), $T_{d,p}$, $T_{d,a}$, $T_{d,h}$ denote the feasible departure time windows (in hours per day) for each vehicle type.
@@ III-B Dynamic Trip Equity
Instead of evaluating the DTE among all travelers in $\mathcal{N}(t)$, which can be computationally inefficient, we propose the DTE defined within the RRCs for each vehicle based on their route plans, ignoring indirect couplings between vehicles. Let $\mathcal{C}_{i,p}(t)$, $\mathcal{C}_{i,a}(t)$, $\mathcal{C}_{i,h}(t)$ be the sets of private, autonomous, and ride-hailing vehicles in $\mathcal{C}_{i}(t)$, i.e., $\mathcal{C}_{i}(t)\!=\!\mathcal{C}_{i,p}(t)\cup{\mathcal{C}_{i,a}(t)}\cup{\mathcal{C}_{i,h}(t)}$. We employ the Gini Coefficient [14], an effective measure of inequity in economics and social sciences, to quantify the DTX distribution among travelers. By definition, the DTE for the traveler in vehicle $i$ and those in the RRCs of vehicle $i$ at $t$ is defined as
@@ III-C Dynamic Trip Equity in Free-Flow Networks
This subsection discusses the DTE in a free-flow road network. In scenarios where the infrastructure is sufficient or the number of vehicles in the system is well below the road capacity, congestion is not a concern in route planning. In these cases, vehicles can freely choose the shortest routes for their trips. Let $DTX_{i,0}$ be the DTX for vehicle $i\!\in\!\mathcal{N}(t)$ traversing a free-flow network. Since $\tau_{i}(t)\!=\!\tau_{\min}^{i}$, by (10), (11b), and (14b), we have
@@
For any vehicle $i\!\in\!\mathcal{N}(t)$ traveling in a free-flow road network on the trip $(o_{i},d_{i})$, the best DTE can be achieved between the traveler in vehicle $i$ and those in its RRCs, denoted as $\mathcal{C}_{i,0}(t)$, if it satisfies
@@
Proposition 1 reveals the conditions for maximizing trip equity in a free-flow road network. These conditions impose constraints on road infrastructure, transportation costs, and vehicle scheduling, which collectively determine the achievable DTX for each vehicle. Note that meeting these conditions ensures that the upper bound of the DTE attainable through route optimization is $1$. In other words, the D-RGS can improve trip equity only through balancing the DTX via reducing traffic congestion. However, it cannot mitigate inequities caused by other factors, such as travel costs or limitations inherent to certain vehicle types.